OpenStudy (anonymous):

Part 1: Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring and why? −b b2 − 4ac 2a Part 2: Create your own quadratic equation. Using complete sentences, and your answer from part 1, determine whether your equation can be solved by factoring.

OpenStudy (anonymous):

@Directrix

OpenStudy (usukidoll):

ahhh you mean the discriminant which is b^2-4ac I love that part

OpenStudy (usukidoll):

if you do that first you can replace the result into the formula and it makes solving a bit easier.

OpenStudy (anonymous):

The answer to Part 1 is the discriminant, b² - 4ac. If the discriminant is negative, the roots are complex numbers. If the discriminant is 0, the roots are -b/2a and -b/2a (repeated roots). You could solve by factoring. If the discriminant is positive, and a perfect square, you can solve by factoring. Part 2 Consider the quadratic x² - 4x - 5. The discriminant is b² - 4ac = (-4)² - 4(1)(-5) = 16 + 20 = 36, which is a perfect square. Hence the quadratic factors are (x - 5)(x + 1)

OpenStudy (usukidoll):

well ok you need a perfect square type of HUMPH! fineeeeeeeeeee

OpenStudy (anonymous):

...?

OpenStudy (usukidoll):

I was here first T____________________________________________T

OpenStudy (anonymous):

It dosent matter if we helped the user to understand but not race who gets the medal first.

OpenStudy (usukidoll):

how come you got a medal and I didn't ? >:/

OpenStudy (anonymous):

A person can only give a medal to a person only.

OpenStudy (usukidoll):

-_-

OpenStudy (usukidoll):

yay!